Learning Goal:

To understand how buffers use reserves of conjugate acid and conjugate base to counteract the effects of acid or base addition on pH.

A buffer is a mixture of a conjugate acid-base pair. In other words, it is a solution that contains a weak acid and its conjugate base, or a weak base and its conjugate acid. For example, an acetic acid buffer consists of acetic acid, CH3COOH, and its conjugate base, the acetate ion CH3COO−. Because ions cannot simply be added to a solution, the conjugate base is added in a salt form (e.g., sodium acetate NaCH3COO).

Buffers work because the conjugate acid-base pair work together to neutralize the addition of H+ or OH− ions. Thus, for example, if H+ ions are added to the acetate buffer described above, they will be largely removed from solution by the reaction of H+with the conjugate base:

H++CH3COO−→CH3COOH

Similarly, any added OH− ions will be neutralized by a reaction with the conjugate acid:

OH−+CH3COOH→CH3COO−+H2O

This buffer system is described by the Henderson-Hasselbalch equation

pH=pKa+log[conjugate base][conjugate acid]    

A beaker with 1.10×102 mL of an acetic acid buffer with a pH of 5.000 is sitting on a benchtop. The total molarity of acid and conjugate base in this buffer is 0.100 M. A student adds 7.50 mL of a 0.490 M HCl solution to the beaker. How much will the pH change? The pKa of acetic acid is 4.740.

Express your answer numerically to two decimal places. Use a minus ( − ) sign if the pH has decreased.
Answer #1

Sol.

As Total molarity of buffer = 0.1 M

So , [CH3COOH] + [CH3COO-] = 0.1 ( equation 1 )

Initial pH = 5

Also , pKa = 4.74

So , Using Henderson - Hasselbalch equation ,

initial pH = pKa + log ( [CH3COO-] / [CH3COOH] )

5 = 4.74 + log( [CH3COO-] / [CH3COOH] )

log( [CH3COO-] / [CH3COOH] ) = 5 - 4.74 = 0.26

[CH3COO-] / [CH3COOH] = 100.26 = 1.8197

[CH3COO-] = 1.8197 [CH3COOH]  

Putting this in equation number 1 ,

1.8197 [CH3COOH] + [CH3COOH] = 0.1

[CH3COOH] = 0.0355 M

[CH3COO-] = 1.8197 × 0.0355 = 0.0646 M

As the volume of buffer = 1.10 × 102 mL

So , initial millimoles of CH3COOH

= 0.0355 × 1.10 × 102 = 3.905 mmol

and initial millimoles of CH3COO-

= 0.0646 × 1.10 × 102 = 7.106 mmol

Now , millimoles of HCl added = 0.490 × 7.50 = 3.675 mmol

So , after addition of HCl , millimoles of CH3COOH increases and millimoles of CH3COO- decreases

New millimoles of CH3COOH = 3.905 + 3.675 = 7.58 mmol

New millimoles of CH3COO- = 7.106 - 3.675 = 3.431 mmol

So , Using Henderson - Hasselbalch equation ,

New pH = pKa + log ( New millimoles of CH3COO- / New millimoles of CH3COOH )

= 4.74 + log ( 3.431 / 7.58 )

= 4.40

Therefore , change in pH

= New pH - Initial pH

= 4.40 - 5.00

=   - 0.60